Find the change in internal energy of 2 kg of water when heated from 0 to 4 C . The specific heat is 4200 and densities at 0 and 4 are 999.9and 1000 respectively. Atmospheric pressure is 10^5 Pa

Grade:11

1 Answers

Vikas TU
7 years ago
U = F*n*R*T/2
F = 6 for H2O
n = 2000/18
T = 4-0 = 4
put the values and get the internal energy value.